OpenStudy (zubhanwc3):

1. The continuous function f(x) is defined by \[f(x)=\frac{ e ^{(x-1)^{2}} }{ (x-1)^{2} }\] for \[x \neq 1 \] and f(1) = 1. the function f(x) has derivatives of all orders at x=1.

OpenStudy (zubhanwc3):

1. The continuous function f(x) is defined by \[f(x)=\frac{ e ^{(x-1)^{2}} }{ (x-1)^{2} }\] for \[x \neq 1 \] and f(1) = 1. the function f(x) has derivatives of all orders at x=1. a) Write the first four nonzero terms and the general term of the Taylor series for \[e^{(x-1)^2}\] about x=1 b) Use the Taylor series found in part (a) to write about the first four nonzero terms and the general term of the Taylor series for f(x) about x=1. c) Use the ratio test to find the interval of convergence for the Taylor series found in part (b). d)Use the Taylor series for f(x) about x=1 to determine whether the graph of f(x) has any points of inflection.

OpenStudy (zubhanwc3):

we are also told that the maclaurin series for e^x is \[e^x=1+x+\frac{ x^2 }{ 2! }+...\frac{ x^n }{ n! } + ...\]

OpenStudy (zubhanwc3):

why would we do e^(1-x^2) when its e^(x-1)^2

OpenStudy (perl):

Latex typo *

OpenStudy (perl):

$$ \large e^{{(x-1)}^2}=1+{(x-1)}^2+\frac{ {(({x-1})^2})^2 }{ 2! }+...\frac{ {(({x-1})^2})^n }{ n! } + ... $$

OpenStudy (zubhanwc3):

@perl i get that this would be for a, but how would u use this for b, i dont realy understand stuff about general terms.

OpenStudy (perl):

$$ \therefore \text{ means therefore} $$

OpenStudy (perl):

General term is $$ \large {e^{{(x-1)}^2}=1+{(x-1)}^2+\frac{ {(({x-1})^2})^2 }{ 2! }+...\frac{ {(({x-1})^2})^n }{ n! } + ... \\ \iff \\{e^{{(x-1)}^2}={(x-1)}^{2\cdot0}+{(x-1)}^{2\cdot1}+\frac{ {({x-1}})^{2\cdot 2} }{ 2! }+...\frac{ {({x-1})^{2n}} }{ n! } + ...} \\ \therefore \\ e^{{(x-1)}^2}= \sum_{k=0}^{\infty}\frac{(x-1)^{2k}}{k!} } $$

OpenStudy (zubhanwc3):

@perl what is the difference between a and b?

OpenStudy (zubhanwc3):

i mean, how do u incorporate the change in equation? do u make the bottom (x-1)^2 k! ?

OpenStudy (perl):

b is the general term. a) is the first 4

OpenStudy (perl):

err, the first 4 are

OpenStudy (zubhanwc3):

@perl isnt b asking about the first 4 terms of f(x) and \[f(x)=\frac{ e ^{(x-1)^{2}} }{ (x-1)^{2} }\] so how would u go about to solving this?

OpenStudy (perl):

$$ a) \large \text {The first four terms are } \\ \large {e^{{(x-1)}^2}=1+{(x-1)}^2+\frac{ {(({x-1})^2})^2 }{ 2! }+\frac{ {(({x-1})^2})^3 }{ 3! }\\ =1+{(x-1)}^2+\frac{ {({x-1})^4} }{ 2 }+\frac{ ( x-1)^ 6 }{ 6 } \\ \large \text { The general term is }\\ \\{e^{{(x-1)}^2}={(x-1)}^{2\cdot0}+{(x-1)}^{2\cdot1}+\frac{ {({x-1}})^{2\cdot 2} }{ 2! }+...\frac{ {({x-1})^{2n}} }{ n! } + ...} \\ \therefore \\ \\ e^{{(x-1)}^2}= \sum_{k=0}^{\infty}\frac{(x-1)^{2k}}{k!} } $$

OpenStudy (perl):

b) $$ \large { e^{{(x-1)}^2}= \sum_{k=0}^{\infty}\frac{(x-1)^{2k}}{k!} \\ \therefore \\ \Large \frac{e^{(x-1)^2}}{(x-1)^2} = \frac{\sum_{k=0}^{\infty}\frac{(x-1)^{2k} }{k!} }{(x-1)^2} = \sum_{k=0}^{\infty}\frac{(x-1)^{2k} }{k!(x-1)^2} \\\text{ }\\ \Large =\sum_{k=0}^{\infty}\frac{(x-1)^{2k-2} }{k!} } $$

OpenStudy (zubhanwc3):

and what steps do u take for c? i have a general idea of how to do the ratio test, but im not too sure

OpenStudy (zubhanwc3):

so i would get (x-1)^2 / (n+1) i would make the abs value of that = 1? or would imake it equal to 0, if 0, than y?

OpenStudy (perl):

the limit of that is zero

OpenStudy (zubhanwc3):

why would the limit be 0?

OpenStudy (perl):

$$ \\\Large { \lim_{n \to \infty} |\frac{a_{n+1}}{a_n}| \\[0.2 in] \\ \\ \lim_{n \to \infty} |\frac{ \frac{ (x-1)^{2(n+1)-2} }{(n+1)!}} {\frac{(x-1)^{2n-2} }{n!}}| \\ = \lim_{n \to \infty} |{ \frac{ (x-1)^{2n} }{(n+1)!} \cdot \frac{ n! }{(x-1)^{2n-2} } } | \\ = \lim_{n \to \infty} |\frac{(x-1)^2}{n+1}|\\ = (x-1)^2 \cdot \lim_{n \to \infty} | \frac{1}{n+1}| = (x-1)^2 \cdot 0 = 0 } $$

OpenStudy (perl):

Since the limit is always less than 1, the series converges for all x.

OpenStudy (zubhanwc3):

so the interval of convergance would be 0?

OpenStudy (perl):

the interval of convergence is all real numbers

OpenStudy (zubhanwc3):

and how would u solve d?

OpenStudy (perl):

you have a mistake , the maclaurin series for e^x is about x = 0, not x = 1

OpenStudy (zubhanwc3):

no, he gives us that maclaurin series at the start, than the questions are about the taylor series where x = 1

OpenStudy (perl):

Write the first four nonzero terms and the general term of the Taylor series for $$ e^{(x-1)^2}$$ about x=1. We plugged (x-1)^2 into the maclaurin series for e^x, which is centered at x = 0. we cant do that

OpenStudy (zubhanwc3):

but if u center it at 1, x-1 is 0, which means that the graph would only shift, right?

OpenStudy (perl):

hmm, we might be able to do that

OpenStudy (zubhanwc3):

so how would we do d?

OpenStudy (zubhanwc3):

i know that inflection is second derivative so do i find the second derivative and plug it in?